OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would multiply the following expression: ^12√x^5 times ^6√x (2 points) Part 2: Write the simplified answer in radical form. (2 points)

OpenStudy (anonymous):

Hint... \[\huge \sqrt[12]{x^5}\cdot \sqrt[6]{x}=\left(x\right)^{\frac{5}{12}}\cdot \left(x\right)^{\frac16}\] now use this formula... \[\huge x^a\cdot x^b=x^{a+b}\]